# Cluster and Predict Latent Patches for Improved Masked Image Modeling

### v2

@@ 4.2 Ablation Studies
Predictor architecture. We evaluate the different predictor architectures. The split predictor produces better representations while training 32% faster than the fused predictor (table 1(a)). Using pure cross-attention in the predictor obtains better representations and allows an additional 18% speedup by avoiding a forward on patch tokens. It also removes the dependency between different predictions, alleviating the need for repeated forwards of the predictor as in I-JEPA.
@@
Loss formulation. We evaluate the different strategies for computing a loss function discussed in Sect. 3.1. We compare the performance of direct Huber loss, an iBOT loss with an MLP head, as well as our clustering-based loss with a linear or MLP head. With no head and direct loss, the model starts well but quickly regresses to worse representations. The iBOT head does not work in our setup, probably because of the split predictor design. Finally, the proposed clustering head alleviates all these issues and allows for stable training and good representations.
@@
Registers. In our model, the local feature maps serve as a supervisory signal to train the model, so high-quality feature maps are crucial. To this end, we use register tokens, which were proposed to improve the quality of feature maps. We see that using registers has a large effect on performance, with an improvement of $3.2$ points on ADE and $2.1$ points on ImageNet when using 16 registers (table 1(g)).
@@ 4.3 Results
Model Arch. Dataset val v2 ReaL A ObjectNet iNat Places205 SUN397
@@
Table 2: Image classification results. For each baseline method, we report the model size closest to ViT-L/14.
Image classification. We evaluate our model and compare it to state-of-the-art reconstruction-based SSL models. We run the evaluation on four datasets including object recognition, fine-grained classification, and scene recognition. We use ImageNet-1k (Russakovsky et al., 2015), iNaturalist 2021 (Van Horn et al., 2021), Places205 (Zhou et al., 2017), and SUN397 (Xiao et al., 2010). For ImageNet, we also report OOD robustness by running inference on additional test sets: ImageNet-V2 (Recht et al., 2019), ImageNet-ReaL (Beyer et al., 2020), ImageNet-A (Hendrycks et al., 2021), and ObjectNet (Barbu et al., 2019). For each model, we resize the image to 224$\times$224 and collect the patch tokens output by the model. We feed those to an attentive probe implemented as a single layer of multi-head cross-attention with a single query (head size $d//64$, no residual). For $c$ classes and an embedding size $d$, the probe contains $2d^{2}+(3+c)d$ parameters, which are trained with AdamW for $10$ epochs, selecting the best learning rate for each model/task on a held-out split of the training set. More details on the protocol are available in appendix F. All the results are summarized in Table 4.3.
We see that our model outperforms all previous state-of-the-art models, by a large margin. When training on ImageNet-1k, we observe very good results, outperforming all other reconstruction-based models of comparable size. CAPI excels particularly on out-of-distribution generalization, outperforming all baselines by more than 19 points on ObjectNet. Additionally, while the gap with other methods is somewhat limited on ImageNet, the difference in scene classification (SUN397) is much larger. Interestingly, we observe that CAPI works particularly well on larger and more diverse datasets. Our three CAPI models outperform all baselines on all datasets, except the CAPI-Places205 which is slightly below AIM on ImageNet-A. It should be noted, however, that AIM was trained on more than 2 billion images, with a sampling distribution tailored towards ImageNet. CAPI significantly reduces the gap between reconstruction-based methods and our topline DINOv2+reg: the gap on ImageNet goes from 8.4 points to 3.6, and on SUN397 goes from 13.2 to 1.8.
@@
Table 3: Comparison with the state of the art on image segmentation using frozen features. We report both $k$-NN and linear segmentation performance. For reference, we also report the performance of some other non-MIM SSL models. This shows that CAPI narrows the gap using only a MIM approach.
Dense image understanding. As we have seen above, our model allows training high-quality local features that can be successfully pooled to solve image-level tasks. We also want to evaluate our model on dense prediction problems such as image segmentation. To this end, we run $k$-NN and linear segmentation following the protocol described in the experimental details. We run this for all the baselines reported above on three datasets. We use ADE-20k (Zhou et al., 2017), Pascal VOC 2012 (Everingham et al., 2010), and Cityscapes (Cordts et al., 2016). We report mIoU for all configurations in Table 4.3.
As in the classification evals, CAPI trained on ImageNet-1k outperforms all reconstruction-based baselines by quite a wide margin on all evaluation setups. When training on larger datasets, the conclusion is similar: CAPI outperforms all baselines by a wide margin and even beats DINOv2+reg in some setups. When trained on Places205, CAPI achieves mIoU 1.2 points higher than DINOv2+reg. To our knowledge, this is the first time DINOv2+reg is bested on segmentation with frozen features on ADE20K. DINOv2+reg, however, remains the most versatile model, with good results across the board and the best results on Cityscapes.
@@ 4.4 Additional explorations
High-resolution image understanding. Our model was trained on $224\times 224$ images. To compare with the I-JEPA model trained natively at high resolution, we try to evaluate our model on $448\times{}448$ images. In table 4, we see that our model does not require high-resolution training or evaluation to achieve the best performance. Our model trained at 224 and evaluated at 224 outperforms the large I-JEPA model trained at 448 and evaluated at 448. Moreover, using RoPE embeddings, CAPI is more robust to resolution changes: it loses only $0.5\%$ (versus $1.0\%$ for I-JEPA) when increasing the resolution.
@@ 5 Discussion and Concluding Remarks
In this paper, we have proposed a novel reconstruction-based self-supervised learning algorithm. Our algorithm is based on an online clustering of dense features computed with a teacher network. The latent assignments are used as targets to train the student. We propose to implement the student as an encoder followed by a predictor: a cross-attention decoder. The teacher is updated as an EMA of the encoder. The proposed algorithm is simple and allows the training of a state-of-the-art model. Our ViT-L outperforms all available reconstruction-based models, including much larger architectures. We have shown promising scaling trends until the 300M model sizes of the ViT family, opening up a potential for further scaling in future work.
@@ Acknowledgments
Appendix A Detailed overview
In Figure 8, we provide a detailed overview of the complete method, with tensor sizes annotated for a reference CAPI ViT-L/14 model.
Figure 8: Detailed overview of our method with reference tensor sizes for a CAPI ViT-L/14 model. We denote in red the parts that are trained by the main loss, in purple the parts that are trained with the clustering loss, and in blue the parts that are updated by the EMA.
Appendix B Loss curve
We report in Figure 10 the loss curve of our CAPI ViT-L model. After an initial adjustment period, the loss trends smoothly downwards, with no sign of instability or plateauing. Compared to other latent masked image modeling methods such as I-JEPA or iBOT, this trend is reassuring, and might indicate good potential for further scaling.
@@ Appendix B Loss curve
Appendix C Blockwise masking strategy
@@ Appendix D Self-distillation interpretation
It was observed in DINO (Caron et al., 2021) that the downstream scores of the EMA model were consistently higher than the ones of the online model during training. This led to the interpretation of DINO as a self-distillation method, where the EMA model, the "teacher" distilled its slightly better representations into the online model, the "student". We observe that this interpretation still seems to hold in CAPI, albeit to a lesser extent, as evidenced by the comparison of teacher and student performance in Figure 10.
Appendix E Modified Sinkhorn-Knopp
We provide the pseudo-code for the standard Sinkhorn-Knopp and for our modified version in Figure 11. Both the original code and the proposed change are very simple. The actual code additionally contains an initial additive shift to prevent numerical instabilities in the exponential, as well as a collective all_reduce for distributed training.
@@ Appendix E Modified Sinkhorn-Knopp
Appendix F Detailed evaluation protocol
@@ F.4 Baselines
Appendix G Compute cost and environmental footprint
@@
Appendix H List of models used
@@
Appendix I Visualisations
In Figures Figure 12 and 13, we provide visualisations of the feature maps of CAPI compared to other state-of-the-art self-supervised vision models.

### v3

@@ 4.2 Ablation Studies
Predictor architecture. We evaluate the different predictor architectures. The split self-attention predictor produces better representations while training 32% faster than the fused predictor (table 1(a)). Using pure cross-attention in the predictor obtains even better representations and allows an additional 18% speedup by avoiding a forward on patch tokens. It also removes the dependency between different predictions, alleviating the need for repeated forwards of the predictor as in I-JEPA.
@@
Loss formulation. We evaluate the different strategies for computing a loss function discussed in section 3.1. We compare the performance of direct Huber loss, an iBOT loss with an MLP head, as well as our clustering-based loss with a linear or MLP head. With no head and direct loss, the model starts well but quickly regresses to worse representations. The iBOT head does not work in our setup, probably because of the split predictor design (since it implies that target tokens and prediction tokens do not come from the same network). Finally, the proposed clustering head alleviates all these issues and allows for stable training and good representations.
@@
Registers. In our model, the local feature maps serve as a supervisory signal to train the model, so high-quality feature maps are crucial. To this end, we use register tokens, which were proposed to improve the quality of feature maps. We see that using registers has a large effect on performance, with an improvement of $3.2$ points on ADE and $2.1$ points on ImageNet when using 16 registers (table 1(g)). This is coherent with recent observations on tabular data, for which registers were crucial to learn good representations. (Thimonier et al., 2024).
@@ 4.3 Results
Model Arch. Dataset val ReaL v2 A ObjectNet iNat Places205 SUN397
[…]
@@
[…]
Image classification. We evaluate our model and compare it to state-of-the-art reconstruction-based SSL models. We run the evaluation on four datasets including object recognition, fine-grained classification, and scene recognition. We use ImageNet-1k (Russakovsky et al., 2015), iNaturalist 2021 (Van Horn et al., 2021), Places205 (Zhou et al., 2017), and SUN397 (Xiao et al., 2010). For ImageNet, we also report OOD robustness by running inference on additional test sets: ImageNet-V2 (Recht et al., 2019), ImageNet-ReaL (Beyer et al., 2020), ImageNet-A (Hendrycks et al., 2021), and ObjectNet (Barbu et al., 2019). For each model, we resize the image to 224$\times$224 and collect the patch tokens output by the model. We feed those to an attentive probe implemented as a single layer of multi-head cross-attention with a single query (head size $d//64$, no residual). For $c$ classes and an embedding size $d$, the probe contains $2d^{2}+(3+c)d$ parameters, which are trained with AdamW for $10$ epochs, selecting the best learning rate for each model/task on a held-out split of the training set. More details on the protocol are available in appendix H. All the results are summarized in section 4.3.
We see that our model outperforms all previous state-of-the-art models, by a large margin. When training on ImageNet-1k, we observe very good results, outperforming all other reconstruction-based models of comparable size. CAPI excels particularly on out-of-distribution generalization, outperforming all baselines by more than 19 points on ObjectNet. Additionally, while the gap with other methods is somewhat limited on ImageNet, the difference in scene classification (SUN397) is much larger. Compared to iBOT, which uses a DINO loss to stabilize a MIM loss, CAPI performs better across the board. MIM-refiner, which uses a contrastive loss to refine a pretrained MIM model, obtains great scores when applied to dBOT. Our CAPI is comparable on most ImageNet benchmarks (+0.1 on IN-val, -0.3 on IN-ReaL, -0.5 on IN-v2) while being slightly better on ImageNet-A (+2.8) and falling behind on ObjectNet (-5.3). CAPI however significantly outperforms MIM-refiner out of domain, with a gap of 3.9, 2.2 and 2.2 points on iNaturalist, Places and SUN respectively.
Interestingly, we observe that CAPI works particularly well when trained on larger and more diverse datasets. In a second category, we compare models trained on datasets larger than ImageNet-1k. Our three CAPI models outperform all baselines on all datasets, except the CAPI-Places205 which is slightly below AIM on ImageNet-A. It should be noted, however, that AIM was trained on more than 2 billion images, with a sampling distribution tailored towards ImageNet. When trained on bigger datasets, CAPI significantly reduces the gap between reconstruction-based methods and our topline DINOv2+reg: the gap on ImageNet goes from 8.4 points to 3.6, and on SUN397 goes from 13.2 to 1.8.
@@
Table 3: Semantic segmentation results. We train a $k$-NN or a logistic regression probe on top of frozen features. For each baseline, we report the model size closest to ViT-L/14. We separate models trained on ImageNet-1k only, and the ones trained on a larger dataset. We report mIoU on all datasets. For reference, we also report the performance of some other SSL models using DINO or contrastive losses to complement the MIM loss. This shows that CAPI narrows the gap using only a MIM approach. : Non-MIM SSL approaches.
Dense image understanding. As we have seen above, our model allows training high-quality local features that can be successfully pooled to solve image-level tasks. We also want to evaluate our model on dense prediction problems such as image segmentation. To this end, we run $k$-NN and linear segmentation following the protocol described in the experimental details. We run this for all the baselines reported above on three datasets. We use ADE-20k (Zhou et al., 2017), Pascal VOC 2012 (Everingham et al., 2010), and Cityscapes (Cordts et al., 2016). We report mIoU for all configurations in section 4.3.
As in the classification evals, CAPI trained on ImageNet-1k outperforms all reconstruction-based baselines by quite a wide margin on all evaluation setups. Compare to dBOT-refined, interestingly, CAPI obtains similar score in linear probing while having worse scores in $k$-nn probing. It is possible that the contrastive refining process is helpful to obtain a good $k$-nn score. When training on larger datasets, the conclusion is similar: CAPI outperforms all baselines by a wide margin and even beats DINOv2+reg in some setups. When trained on Places205, CAPI achieves mIoU 1.2 points higher than DINOv2+reg on ADE20K $k$-nn segmentation. To our knowledge, this is the first time DINOv2+reg is bested on segmentation with frozen features on ADE20K. DINOv2+reg, however, remains the most versatile model, with good results across the board and the best results on Cityscapes.
@@ 4.4 Additional explorations
High-resolution image understanding. Our model was trained on $224\times 224$ images. To compare with the I-JEPA model trained natively at high resolution, we try to evaluate our model on $448\times{}448$ images. In table 4, we see that our model does not require high-resolution training or evaluation to achieve the best performance. Our model trained at 224 and evaluated at 224 outperforms the large I-JEPA model trained at 448 and evaluated at 448. Moreover, using RoPE embeddings, CAPI is more robust to resolution changes: it loses only $0.3\%$ (versus $1.0\%$ for I-JEPA) when increasing the resolution.
@@
5 Broader impact
As with any machine learning model, the impact of our work depends on how it is used. This work in particular focuses on improving the quality of self-supervised visual features, which can pose specific privacy issues if a model is trained on identifiable pictures, and pose fairness issues if the model absorbs biases from the dataset it is trained on. Additionally, the training of our models and the related exploratory experiments directly use energy and computational resources, which can have an environmental impact. We discuss environmental impact in more detail in the appendix.
@@ 6 Discussion and Concluding Remarks
In this paper, we have proposed a novel reconstruction-based self-supervised learning algorithm. Our algorithm is based on an online clustering of dense features computed with a teacher network. The latent assignments are used as targets to train the student. We propose to implement the student as an encoder followed by a cross-attention predictor. The teacher is updated as an EMA of the encoder. The proposed algorithm is simple and allows the training of a state-of-the-art model in the category of reconstruction-based approaches. Our ViT-L outperforms all available reconstruction-based models, including much larger architectures. We have shown promising scaling trends until the 300M model sizes of the ViT family, opening up a potential for further scaling in future work. Some questions remain open: in particular, we believe it would be interesting to explore the scaling capabilities beyond the scale we explored, and to study the optimal training dataset for this sort of training. The question of whether such a pure reconstruction-based approach can reach the performance of contrastive learning methods remains open. We hope that our work will inspire further research in this direction.
@@ Acknowledgments
Appendix A Loss curve
We report in fig. 10 the loss curve of our CAPI ViT-L model. After an initial adjustment period, the loss trends smoothly downwards, with no sign of instability or plateauing. Compared to other latent masked image modeling methods such as I-JEPA or iBOT, this trend is reassuring, and might indicate good potential for further scaling.
@@ Appendix A Loss curve
Appendix B Blockwise masking strategy
@@
Appendix C Batch size
In all our experiments, we use a batch size of 16384. Although the memory footprint is slightly offset by the high proportion of tokens not processed by the encoder and predictor, we recognize that this may be a limiting factor for some users. For the ViT-L, this batch size fits in 4 nodes of 8 A100 80GB GPUs. We believe that the batch size could be reduced to 4096 without significant loss in performance, but the recipe may need adjustments, and we do not have results at smaller batch size.
@@ Appendix D Self-distillation interpretation
It was observed in DINO (Caron et al., 2021) that the downstream scores of the EMA model were consistently higher than the ones of the online model during training. This led to the interpretation of DINO as a self-distillation method, where the EMA model, the "teacher" distilled its slightly better representations into the online model, the "student". We observe that this interpretation still seems to hold in CAPI, albeit to a lesser extent, as evidenced by the comparison of teacher and student performance in fig. 10.
Appendix E Predictor pooling versus attentive probing
We report in appendix E the comparison of linear probing scores on ImageNet using different representations, with the topline being the attentive pooling results. We observe that the predictor learns a much better pooling than the baseline average pooling, although it does not reach the topline where the pooling is supervised (attentive probing).
{tabu}
ll ccccccc IN1k
Representation Learnable head val ReaL V2 A R Sketch Objectnet
avg pooling Linear 76.8% 82.5% 64.4% 21.8% 43.5% 33.9% 33.7%
pred pooling Linear 81.8% 86.9% 71.6% 46.7% 55.8% 47.0% 48.3%
cat(pred pooling, avg pooling) Linear 81.9% 87.0% 71.5% 45.8% 55.6% 46.5% 47.1%
feature map Attentive 83.8% 88.2% 74.8% 56.5% 58.8% 47.9% 54.9%
Table 6: Comparison of linear probing scores on ImageNet using different representations and classification heads. In the first three rows, the pooling is a fixed function, while in the fourth row, the pooling is learned specifically or ImageNet via an attention layer. We observe that, while the average pooling of CAPI does not produce a great global vector, the predictor pooling trick allows CAPI to be competitive on the ImageNet linear probing benchmark. The pooling, however, is not the optimal pooling for ImageNet since an attentive pooling manages to get a better score.
Appendix F PCA visualizations of early models suffering positional from collapse
We show in fig. 11 the PCA visualization of the patch embeddings of an early CAPI model that showed some partial positional collapse.
input PCA 1 PCA 2 PCA 3 PCA 4 PCA 5 PCA 6
Figure 11: PCA visualization of the patch embeddings of some early CAPI models, showing partial positional collapse. The models did learn some useful nontrivial features (around 60% accuracy on ImageNet), but the positional information was very present and dominated the trainingn, preventing the model from learning better features.
Appendix G Modified Sinkhorn-Knopp
We provide the pseudo-code for the standard Sinkhorn-Knopp and for our modified version in fig. 12. Both the original code and the proposed change are very simple. The actual code additionally contains an initial additive shift to prevent numerical instabilities in the exponential, as well as a collective all_reduce for distributed training.
@@ Appendix G Modified Sinkhorn-Knopp
Appendix H Detailed evaluation protocol
@@ H.4 Baselines
Appendix I Compute cost and environmental footprint
@@
Appendix J List of models used
@@
Appendix K Visualisations
In fig. 13 and fig. 14, we provide visualisations of the feature maps of CAPI compared to other state-of-the-art self-supervised vision models.